UT Apr 16 2022 - Split Night LBTB Partner Science 1st half / DEC time 2nd half



Night Info

  • Moonrise 17:52
  • Sunset 18:50
  • Nautical Twilight Ends 19:45
  • Astro Twilight Ends 20:16
  • Night Center 00:19 - Handover from LBTB to LBTO at night center (both sides using LUCIs)
  • Astro Twilight Starts 04:22
  • Nautical Twilight Starts 04:53
  • Sunrise 05:49
  • Moon illumination 0.995

Staff

  • OSA: J.Williams
  • ISp: C.Bolyard/Mtn
  • Mtn.Manager: J.Prothro
  • ISA/Observer: B.Rothberg, R.T.Gatto
  • Special Guests:

Night Plan

Possible Tasks

  • DONE DX ASM Field Aberration Data Collection (use RFBG) - 2hr
    • Obtain data for binodal astrigmaism and alignment of secondary
    • seeing 1" - ideal, seeing up to 1.2" acceptable
    • DX data must be analyzed and confirmation run once completed
    • May wish to collect two sets
    • Potential Fields (variety to account for winds) - /home/lbto/idl/wfsc/field/lists/
      • 1st half (if needed next night)
        • BS9122 07hr +48dec - twilight-22hr
        • BS9126 07hr +29dec - twilight-22hr
        • BS9134 08hr +05dec - twilight-22hr
        • BS 9152 13hr +52dec - all night
      • 2nd half
        • BS9152 13hr +52dec - all night
        • BS9155 14hr -00dec - 22hr-twilight
        • DONE BS9164 16hr +05dec - 23:30-twilight
        • BS9169 17hr +54dec - 22:30 twilight
        • BS9170 17hr -00dec - 00:30-twilight
        • BS9177 18hr +49dec - 00hr-twilight
  • DONE On-sky checks of DX ASM - misc verification that everything works -0.5hr-1hr
  • Other Engineering Tasks
    • MODS spectrophotometric standards - all modes - if clear
      • needed to compare with SX and Rigid secondaries
    • LBC Extra Focal images in equilibrium -IT 8608, IT 8592 - 1hr
      • telescope is in good equilibrium, collimate and then inject Z4/Z11 after which we would take extra-focal (dZ=-0.8) images with offsets to Z22 and Z11
      • seeing < 1.2", any clouds
  • D-time Science
    • V1674 Her Followup - MODS Bino - M.Wagner - 0.75hr
      • Visible 02hr-twilight
      • seeing < 1.8" and clouds 1-2mag
      • OT Files and finder chart: /home/lbto/Service_Observing/2022A_D_E/MODS/MWagner/V1674Her
      • Spectrophotometric Std: BD+33 2642
    • mid-z ULIRG FF0937 (bright enough not too affected by Moon) - 1.5hr
      • seeing < 1.2"
      • thin clouds
    • PHAs - LBCs
      • 2019 XO2 recovery (42, 45, and 52 deg from Moon over weekend)

Poor Weather Backup

Eng Tasks
  • Mixed-mods on-sky LUCI+MODS test -0.5hr
    • send active bino preset
    • GSPC P330E (WD) in Mixed mode folder
    • OT: /home/lbto/Service_Observing/2022A_D_E/LUCI/Restart_Scripts/LUCI_2021_Restart_V1.70.xml
    • R=12.8, K=11.3 - 5pt dither +/-10" for LUCI and MODS - active presets, Guidestars R~11-12
    • Verify closed dome test
  • LUCI Photometric Standards (any seeing, if so bad can't guide, use track mode) - requires clear skies
    • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
    • Select stars based on visibility, need to get stars at two different airmasses
  • Pointing Model - DX needs improvement after ASM re-install
    • Seeing up to ~2.5"
    • Use LUCIs
      • N375 cameras, N30_FieldStop (aka the "Pointing") masks, K-band, 2.51 sec NDIT=2
      • no flexure compensation (doesn't matter with the pointing mask in)
      • Before taking pointing data, take at least 2 images with dROT=180 for rotator center (I usually do 4)
      • use PtStars in IRAF - current or near current LST - re-run every 2-4 stars with current LST
      • active preset to collimate on-axis, park probe, then take image - 1st image
      • turn PCS logging ON (it is possible this is always on now...check)
      • send track mode presets to stars after that Note: if you use off-axis stars and let it collimate at each point, you can also make a collimation model
      • Turn PCS logging off, note filename
Science
  • Rothberg - Local Mergers
    • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V42.xml
    • Not in order of priority
      • NGC 4004 - 21hr-twilight Dec +27
      • NGC 3597 - 23hr-02hr Dec -23
      • NGC 6598 - 03hr-twilight Dec +69
      • UGC 9829 - 01:30hr-twilight Dec -01
      • UGC 10607 - 02hr -twilight Dec +26
      • UGC 10675 - 02hr-twilight Dec +31


Night Log

  • UT 05:28 - clouds in and out all night - seeing has been <1.2" so far
  • UT 06:43 - despite wanting to continue J.Heidt is handing over to LBTO for DEC time
  • UT 06:45 - looking for things to do with R>>10 0
  • idling
  • AllSkyCurrentImage-1.JPG

Field Aberration Collection

  • UT 07:20 - might be clearing will prep for field aberration collection (Seeing so far has been stable)
  • UT 07:22 - J.Williams doing a pointing and collimation check near BS 9164
  • Nice, we're pointing to a hole in the sky
  • UT 07:29 sending active preset to BS9164 - collimate on star to get idea of seeing before proceeding
  • UT 07:32 - send GUIDE preset to BS9164
  • running this on OBS-2 w/another X2GO to avoid SSH/X slowness
  • UT 07:36 - staring field data collection - seeing 1-1.2" on guiders
    • directories are data/20220416/sx_field and dx_field
  • UT 07:56 GS0 on DX needed more time to collimate
  • UT 08:01 - clouds seem to be stabilizing the seeing - getting 0.7"-0.9" on the guiders
  • UT 08:05 - Star 7 failed on SX, will restart at next, Star 20
  • UT 08:08 - on-axis star is now looking faint on DX ... we'll see what we can do
  • UT 08:10 - failed on GS20 - starting on GS8 (skipping G20 and GS7)
  • UT 08:14 - winds reaching towards 20m/s, up to GS10 on SX (after skipping 2) and GS 5 on DX
  • UT 08:32 - DX missed GS20, restarted and accidentally sent it to 10 again, sky cleared out and now its able to doit
  • UT 08:41 - SX completed
  • UT 08:54 - DX still going - somce clouds coming back in from west, seeing has been 1" or better most of the time so far
  • UT 09:01 - DX still struggling with collimation on GS0
  • UT 09:14 - clouds rolling back in - not quite as thick, patchy spots
  • UT 09:29 - DX completed

Backup Science Target Mergers

NGC 6598
  • Decided to go to a backup science as the clouds are starting to thicken up and roll in - will do this target and gauge conditions
  • UT 09:32 - sending preset for Telluric BD+41 300B
  • UT 09:34 - missed the GS on SX
  • UT 09:35 - sending pointing
  • UT 09:37 - sending preset to telluric
  • AllSkyCurrentImage-2.JPG
  • also skirting just under the wind limits
  • Its a double-star but both are actually A1V stars - picked the "B" star based on 2MASS image at PA=0
    • L1 -8.8173", 4.1206"
    • L2 -8.4347", 3.078"
  • UT 09:47 - starting spectra - clouds thickening up and rolling in - heavier bank is to the south, luckily target is to the north
  • UT 09:48 - seeing blew up too
  • UT 09:49 - have to restart RTD on LUCI2
  • UT 09:50 - guidestars are vanishing - they are R=13
  • UT 09:51 - faint scatterd light from companion A0 - its not in the slit - but light is certainly scaterring into the slit
    • L1 0038-0043
    • L2 0039-0044
  • UT 09:52 - sending preset to NGC 6598 - gonna give it the 'ole college try
    • guidestar is R=13.6 for target field
  • AllSkyCurrentImage-3.JPG
    • L1 -0.6774", -0.9730"
    • L2 -0.3118", -1.7255"
    • RTD LUCI-2 is showing RA/Dec coords not X,Y, reset all, STOP , START fixed it - below is the dump
    • Command> Start Aladin executing
      trace 4; grid on; reticle large; setconf sync on; gc off; setconf frame=XY fits; rm -all
      [trace 4]...
      Trace on (level 4)
      [grid on]...
      >>>> Command.exec() : execute "grid on"...
      [reticle large]...
      >>>> Command.exec() : execute "reticle large"...
      [setconf sync on]...
      >>>> Command.exec() : execute "setconf sync on"...
      [gc off]...
      >>>> Command.exec() : execute "gc off"...
      [setconf frame=XY fits]...
      >>>> Command.exec() : execute "setconf frame=XY fits"...
      >>>> Command.execSetconfCmd(frame=XY fits) => prop=[frame] value=[XY fits])
      [rm -all]...
      >>>> Command.exec() : execute "rm -all"...
      Start Aladin done
      Aladin accessed from thread RMI TCP Connection(74)-192.168.60.12
      Aladin idle. Queueing Load file /home/lbto/.luci//readout2/luci2.20220416.0048.fits
      Width:64 Bottom:1920 Below:1856 Top:1983 Above:2047 Rows: 64 128
      Pixels: 968/1726, multiplier: 2.45 Rows changed: 64 from 1920 to 1983
      *** stream => detect: FITS XFITS
      Wait for Aladin plane luci2.20220416.0048 executing
      sync
      *** MultiExtension 0 detect => FITS XFITS
      . Creating the Image plane luci2.20220416.0048[0]...
      -- Loading FITS image
      [sync]...
      >>>> Command.exec() : execute "sync "...
      *** => NAXIS1=2048 NAXIS2=2048 BITPIX=32 => size=16777216
      *** => BZERO = 0.0
      *** => BZERO = 0.0 BSCALE = 1.0
      *** => Reading and autocutting in 0.053 s => 301.89 Mbyte/s
      *** Original pixels saved in cache [/home/lbto/.aladin/.aladin.64950/f24.tmp]
      *** Original pixels RAM free for luci2.20220416.0048[0]
      *** Plane [luci2.20220416.0048] load&build in 66ms
      *** Checking language support...
      *** Add [http://aladin.u-strasbg.fr/java/nph-aladin.pl?frame=getLang] for cache upgrade
      *** [http://aladin.u-strasbg.fr/java/nph-aladin.pl?frame=getLang] read in cache !
      Display FITS file luci2.20220416.0048 executing
      cview luci2.20220416.0048; cm auto BB noreverse
      [cview luci2.20220416.0048]...
      >>>> Command.exec() : execute "cview luci2.20220416.0048"...
      [cm auto BB noreverse]...
      >>>> Command.exec() : execute "cm auto BB noreverse"...
      Display FITS file luci2.20220416.0048 done
      Wait for Aladin plane luci2.20220416.0048 done
      Remove MOS planes executing
      rm CTR_refslits; rm CTR_reftargets; rm LMS_refslits; rm LMS_reftargets
      [rm CTR_refslits]...
      >>>> Command.exec() : execute "rm CTR_refslits"...
      [rm CTR_reftargets]...
      >>>> Command.exec() : execute "rm CTR_reftargets"...
      [rm LMS_refslits]...
      >>>> Command.exec() : execute "rm LMS_refslits"...
      [rm LMS_reftargets]...
      >>>> Command.exec() : execute "rm LMS_reftargets"...
      Remove MOS planes done
      FITS header keyword MRSLITNO not found in luci2.20220416.0048
      SelectedTwo: null
      SelectedTwo: null
      *** Original pixels removed from cache [/home/lbto/.aladin/.aladin.64950/f21.tmp]
      *** Original pixels removed from cache [/home/lbto/.aladin/.aladin.64950/f20.tmp]
  • UT 10:10 - starting science - looks like its in the slit - not 100% certain
  • UT 10:14 - guidestar fading in and out, lets see if we make the first offset...
  • UT 10:18 - not sure I see something in the first dither position - we'll see after 2nd exposure and a pair subtraction
  • UT 10:21 - nothing in guider - have it paused after 2nd exposure - will see if anything appears - and wait
  • SHOCKED! i see a faint trace on both sides - clearly need more S/N - but that's life
  • UT 10:23 - waiting for guidestar to return
  • AllSkyCurrentImage-4.JPG
  • UT 10:33 - guidestar came back - continuing script
  • UT 10:43 - hole opened up in the NW part of the sky which we are aimed at, the rest is still socked in - i'll take what i can get
  • UT 10:49 - closing back up....(sky not telescope)
  • Pausing after 0052 and 0053 - to see if guider remains visible
  • UT 10:52 - continuing to next dither
  • UT 10:58 - GS fading again - pause in place
  • UT 11:05 - still waiting
  • UT 11:12 = putting LUCI into safe mode - closing up
    • L1 0044-0053
    • L2 0045-0054

MISC

  • UT 08:44 - waking MODs just in case it stays good - some programs that can be done
  • UT 09:08 - checked MWagner script - acquisition still was incorrect, used a template and copied and pasted targ info and updated for exptime and LS choice - now looks OK
  • UT 11:23 - put MODS to sleep

Calibrations

Bino-LUCI 0.5" slit - G210 K 2.30microns

* L1 0054-0071
  • L2 0055-0072

  • UT 11:44 - putting LUCI-2 into safe config
  • UT 11:49 - putting LUCI-1 into safe config

-- BarryRothberg - 13 Apr 2022
Topic revision: r10 - 16 Apr 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback